[figure: 9.7 cm]

MAGNETORESISTANCE OF THIN MAGNETIC FILMS WITH

BLOCH-WALL-TYPE MAGNETIC STRUCTURES

K. Mibu^{a}, T. Nagahama^{a}, T. Ono^{b}, and T. Shinjo^{a}

aInstitute for Chemical Research, Kyoto University, Uji, Kyoto-fu 611-0011, Japan

bFaculty of Science & Technology, Keio University, 3-14-1 Hiyoshi, Kohoku-ku, Yokohama 223-8522, Japan

The magnetoresistance originating from magnetic structures with gradually rotating magnetic moments, like a Bloch wall, was investigated using soft-magnetic (NiFe)/hardmagnetic (CoSm) bilayers, which provide a system with well characterized magnetic structures [1]. The magnetoresistance was measured with an electric current parallel and perpendicular to the film plane; the former geometry corresponds to the magnetoresistance measurement with a current parallel to a Bloch wall and the latter corresponds to that with a current perpendicular to a Bloch wall. The observed magnetoresistance curves were explained by combination of anisotropic magnetoresistance, GMR-type magnetoresistance, and some additional magnetoresistance effects.

[1] K. Mibu et al., Phys. Rev. B 58 (1998) 6442.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : K. Mibu

Address for correspondence : Institute for Chemical Research, Kyoto University Uji, Kyoto-fu 611-0011, Japan

Email address : mibu@scl.kyoto-u.ac.jp